OpenStudy (anonymous):

At 20 degrees Celsius the vapor pressure of benzene (C6H6) is 75 torr, and that of toluene (C7H8) is 22 torr. Assume that benzene and toluene form an ideal solution. a) What is the composition in mole fractions of a solution that has a vapor pressure of 45 torr at 20 degrees celsius. b) What is the mole fraction of benzene in the vapor above the solution described in part (a)? Why is the mole fraction of Toluene 1-x?

OpenStudy (anonymous):

lies and slander! (slander is also known as slendermen)

OpenStudy (bahrom7893):

Useless responses are useless

OpenStudy (bahrom7893):

@dumbcow can u help dall out?

OpenStudy (bahrom7893):

I would if i knew chem.. sorry

OpenStudy (anonymous):

you are a great example of one bahrom.

OpenStudy (abb0t):

OHHHHHHHHHHH!!

OpenStudy (anonymous):

now could we get back to the slenderman dallbencaer?

OpenStudy (abb0t):

I don't think this is using the ideal gas law, instead using roultz law: \(P_{tot}=x_ap_a+x_b+p_b...\) Knowing that x_a+x_b = 1 they have all the information, so they should be able to do this. for part b, they can find the vapor presure from benzene using the same formula. find the mole fraction by dividing by total pressure.

OpenStudy (anonymous):

Raoult's Law states that for an ideal solution the partial vapor pressure of a component in solution is equal to the mole fraction of that component times its vapor pressure when pure ok, let's make the mole fraction of benzene be X therefore, the mole fraction of toluene must be 1-X 75X + 22(1-X) = 37 75X + 22 - 22X = 37 53X = 15 X = 15/53 for benzene = 0.28 for toluene the mole fraction would be (53-15)/53 or 38/53 = 0.72

OpenStudy (nincompoop):

ya I should read it again

OpenStudy (nincompoop):

thanks

OpenStudy (abb0t):

where did you get your numbers from @igotreported

OpenStudy (anonymous):

(a) What is the composition in mole fractions of a solution that has a vapor pressure of 40. torr at 20°C? Let X be the mole fraction of benzene in this ideal solution. Thus 1-X is the mole fraction of toluene in this ideal solution. The following is given: 75X + 22(1-X) = 40, or: 53X = 18, and X = 18/53. Benzene's mole fraction is 18/53, and toluene's mole fraction is 35/53. (b) What is the mole fraction of benzene in the vapor above the solution described in part (a)? (75*18/53)/40 = 63.7%

OpenStudy (anonymous):

brain

OpenStudy (abb0t):

LIES AND SLANDER!!!!!!!!!!!

OpenStudy (anonymous):

lies and slendermen abb!

OpenStudy (abb0t):

I don't see where 37 came from!! This is all lies and slander!!

OpenStudy (anonymous):

work the problem

OpenStudy (nincompoop):

let me read on raoult's law brb before I get slandered

OpenStudy (anonymous):

somethings happening to my eyes when i lokk at your profile pic nina

OpenStudy (anonymous):

i think your profile pic has a curse, no joking

OpenStudy (nincompoop):

okay I am done reading it

OpenStudy (anonymous):

and?

OpenStudy (anonymous):

i kinda hurried on the my answer so im not surprised if im wrong

OpenStudy (nincompoop):

I think you missed some big parts

OpenStudy (nincompoop):

let me read your solution first

OpenStudy (anonymous):

okay whats your answer

OpenStudy (nincompoop):

I am curious @igotreported where did you get 37?

OpenStudy (anonymous):

i showed steps... you guys should look at the steps first.

OpenStudy (abb0t):

yeah, that's what I'm wondering wehre the 37 came from. Lol

OpenStudy (abb0t):

Shouldn't it be 45. Lol

OpenStudy (anonymous):

okay thats not funny.... i dont understand why you wrote lol....

OpenStudy (abb0t):

lol = losers occasionally leave

OpenStudy (anonymous):

interesting.. except for i didnt see myself a loser occasionally leaving.... please explain

OpenStudy (anonymous):

oh and btw,, thanks for the medal:)

OpenStudy (anonymous):

LMAO I am still lost as to how you all know that the mole fraction of toluene = 1-X? If Ptot = (Pa)(Xa)+(Pb)(Xb) Ptot = 45 Pa = 75 Xa = mol benzene/mol benzene+mol toluene Pb = 22 Xb = mol toluene/mol benzene+mol toluene thus 45 = (75)(a/a+b)+(22)(b/a+b) and i got nothing from here where is 1 coming from? Lies and Slander I say =-P Night peeps, thanks for helping

OpenStudy (anonymous):

btw if I was a teacher, I would totally know how to do this =-P

OpenStudy (anonymous):

your right